OpenStudy (cherie_magee):

Find the quotient. (4ab - 8a 2 + 12ac) ÷ 4a

OpenStudy (ivettef365):

if you simplify this way is easier 4ab -8a^2 +12ac ----- ------ ------ (now you can simplify) 4a 4a 4a

OpenStudy (cherie_magee):

not exactly @ivettef365

OpenStudy (ivettef365):

ok you can eliminate 4a/4a on first one you have left only b -8/4 = -2 and a^2/a = a, so you have -2a then 12/4 = 3 and eliminate the a because a/a = 1 so you have 3c now combine and you have b-2a+3c